# Holistically-Nested Structure-Aware Graph Neural Network for Road Extraction

### v1

@@ 2 Related Work
The past decade has witnessed the remarkable success of deep learning in various challenging computer vision [70, 70] and remote sensing [33, 53, 13, 15, 1, 46, 16, 17] tasks. One of the first deep learning based road extraction methods was developed by Mnih and Hinton [25], which proposed to adopt restricted Boltzmann machines (RBMs) to detect road from high resolution aerial images. Saito et al. [30] applied Convolutional Neural Network (CNN) to extract both buildings and roads from satellite images without pre-processing step required by [25]. A U-shaped fully convolutional network (FCN) [21] architecture was proposed by Ramesh et al. [18]. This architecture comprised a stack of convolutions followed by corresponding stack of mirrored deconvolutions with the usage of skip connections in between for preserving the local information. Zhang et al. [67] combined the strengths of U-Net [29] and residual learning [11] in training deep neural networks for road extraction. Methods in tackling the small receptive field issue have also been proposed. For instance, Zhou et al. [69] built upon the dilated convolution [4] layers to adjust receptive fields of feature points without decreasing the resolution of feature maps. A comprehensive review of road extraction methods can be found in [14].
@@
Graphical models and graph neural networks (GNNs) have emerged as powerful tools in the field of computer vision [52, 42, 28, 44, 60, 71, 57, 45, 66], offering a versatile framework for representing and processing complex visual data. These approaches leverage the inherent structure and relationships within images [58, 56, 41, 54, 63, 10] and videos [50, 48, 43, 5], allowing for more nuanced and context-aware analysis. By representing visual elements as nodes and their interactions as edges, graphical models and GNNs can capture spatial, temporal, and semantic dependencies that are crucial for tasks such as visual information retrieval [12], stylization [49, 51], object detection [40, 68], scene understanding [7], and image or video segmentation [59, 47, 22, 39]. In recent years, these techniques have also found significant applications in remote sensing and geospatial data analysis.

### v2

@@ 2 Related Work
The past decade has witnessed the remarkable success of deep learning in various challenging computer vision and remote sensing [57, 35, 74, 73, 33, 55, 53, 13, 15, 1, 46, 16, 17] tasks. One of the first deep learning based road extraction methods was developed by Mnih and Hinton [25], which proposed to adopt restricted Boltzmann machines (RBMs) to detect road from high resolution aerial images. Saito et al. [30] applied Convolutional Neural Network (CNN) to extract both buildings and roads from satellite images without pre-processing step required by [25]. A U-shaped fully convolutional network (FCN) [21] architecture was proposed by Ramesh et al. [18]. This architecture comprised a stack of convolutions followed by corresponding stack of mirrored deconvolutions with the usage of skip connections in between for preserving the local information. Zhang et al. [70] combined the strengths of U-Net [29] and residual learning [11] in training deep neural networks for road extraction. Methods in tackling the small receptive field issue have also been proposed. For instance, Zhou et al. [72] built upon the dilated convolution [4] layers to adjust receptive fields of feature points without decreasing the resolution of feature maps. A comprehensive review of road extraction methods can be found in [14].
@@
Graphical models and graph neural networks (GNNs) have emerged as powerful tools in the field of computer vision [54, 42, 28, 44, 63, 75, 60, 45, 69], offering a versatile framework for representing and processing complex visual data. These approaches leverage the inherent structure and relationships within images [61, 59, 41, 56, 66, 10] and videos [50, 48, 43, 5], allowing for more nuanced and context-aware analysis. By representing visual elements as nodes and their interactions as edges, graphical models and GNNs can capture spatial, temporal, and semantic dependencies that are crucial for tasks such as visual information retrieval [12], stylization [52, 49, 51], object detection [40, 71], scene understanding [7], and image or video segmentation [62, 47, 22, 39]. In recent years, these techniques have also found significant applications in remote sensing and geospatial data analysis.